OpenStudy (anonymous):

How can x + 8 = -x + 7 be set up as a system of equations?

OpenStudy (anonymous):

y = x + 8 y = -x + 7 y = x + 8 y = -3x + 7 y = x + 5 y = -2x + 11 y = 8 y = -3x + 7